OpenStudy (anonymous):

how can i determine if graphs are parallel y+8=-6,-2x+y=5

OpenStudy (♪chibiterasu):

To prove two lines are parallel, they must have the same slope.

OpenStudy (♪chibiterasu):

y + 8 - 8 = -6 - 8 y = -14 -2x/-2 + y = 5/-2 x + y - x= -2.5 - x y = -2.5 - x They are not parallel.

OpenStudy (anonymous):

thank you so much.....I have been teaching myself and still question how i get my answer that a lot of the times are correct

OpenStudy (♪chibiterasu):

y = mx + b Remember that m is the slope, that will do you a lot just remembering that. :P

OpenStudy (anonymous):

thanks